The December 31, 2015, balance sheet of Schism, Inc., showed long-term debt of $1,460,000, $152,000 in the stock account and $2,770,000 in the additional paid-in surplus account. The December 31, 2016, balance sheet showed long-term debt of $1,700,000, $162,000 in the common stock account and $3,070,000 in the additional paid-in surplus account. The 2016 income statement showed an interest expense of $100,000 and the company paid out $157,000 in cash dividends during 2016. The firms net capital spending for 2016 was $1,080,000, and the firm reduced its net working capital investment by $137,000. What was the firms cash flow to creditors during 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) Cash flow to creditors What was the firms cash flow to stockholders during 20167 (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) 2 Cash flow to stockholders What was the firms cash flow from assets during 2016? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g.. 1,234,567.) Cash flow from assets What was the firms operating cash flow during 2016? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) Operating cash flow s
Answer #1

1. Cash Flow to creditors = Interest - Increase in Debt = 100,000 - (1,700,000-1,460,000) = -140,000

2.Cash Flow to Equity Holders = Dividends - Increase in Equity = 157000 -((3,070,000+162,000)-(2,770,000+152,000)) = -153,000

3. Cash Flow Assets = Cash Flow to Creditors + Cash Flow to Equity = -140,000 -153,000 = -293,000

4. Cash Flow Operations = Cash flow Assets - Increase in Fixed Assets + Decrease in NWC = -293,000 -1,080,000 + 137,000 =
-1,236,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well